OpenStudy (anonymous):

Solve the system. y = 2 - x and y = 4x + 7 Answer (1, 1) (1, 3) (-1, 3) (-1, 1) no solution

OpenStudy (anonymous):

compare the equations and get the value for x put that value of x in any of the two given equations ad get value for y and get your solution.

OpenStudy (texaschic101):

y = 2 - x now sub 2 - x in for y in the other equation y = 4x + 7 2 - x = 4x + 7 2 - 7 = 4x + x -5 = 5x -5/5 = x -1 = x now sub -1 in for x in either of the equations to find y y = 2 - x y = 2 - (-1) y = 2 + 1 y = 3 now lets check it.... y = 4x + 7 3 = 4(-1) + 7 3 = -4 + 7 3 = 3 (correct) ANSWER : x = -1 and y = 3 or (-1,3)